OpenStudy (anonymous):

Solve the equation for x. Round your answer to two decimal places. Do not include "x =" in your answer. 4^x = 20 plz i need this last answer. my life depends on it

OpenStudy (bahrom7893):

xln4 = ln20

OpenStudy (bahrom7893):

x = ln20/ln4 = 2.16

OpenStudy (anonymous):

YES!!! thank you

OpenStudy (llort):

\[x = (2 i \pi n+\log(5)+2 \log(2))/(2 \log(2)), n \in Z\]

OpenStudy (llort):

I think...

OpenStudy (saifoo.khan):

sorry, idk.

OpenStudy (anonymous):

im in a great mood so u all get medals :)

OpenStudy (amistre64):

kidnappers are holding your little puppy for ransom and demand you to solve this unsolvable problem ... hmmm; tight spot :)

OpenStudy (anonymous):

lol, tho if it was a pupy i would just go in there guns blazing and get my puppy bak

OpenStudy (saifoo.khan):

Oh Cool. do u want one too?

OpenStudy (anonymous):

sure. that would be great

OpenStudy (saifoo.khan):

lol, Given!

OpenStudy (anonymous):

i thank thee very much saifoo

OpenStudy (saifoo.khan):

\begin{array}l\color{red}{\text{Y}}\color{orange}{\text{o}}\color{#9c9a2e}{\text{u}}\color{green}{\text{ }}\color{blue}{\text{w}}\color{purple}{\text{e}}\color{purple}{\text{l}}\color{red}{\text{c}}\color{orange}{\text{o}}\color{#9c9a2e}{\text{m}}\color{green}{\text{e}}\color{blue}{\text{!}}\color{purple}{\text{ }}\color{purple}{\text{:}}\color{red}{\text{)}}\color{orange}{\text{}}\end{array}

OpenStudy (anonymous):

pretty colors